OpenStudy (anonymous):

Which property would allow you to use mental computation to simplify the problem 27 + 15 + 3 + 5?

OpenStudy (anonymous):

wouldn't it be easier to do this in your mind if the 27 and 3 were next to each other, and if the 15 and 5 were together?

OpenStudy (anonymous):

Commutative and Associative Properties of Addition

OpenStudy (aripotta):

associative property of addition

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

I think just commutative...

OpenStudy (anonymous):

\[27+15+3+5=27+3+5+15\] by the commutative law

OpenStudy (aripotta):

well, you could use both

OpenStudy (anonymous):

associative would mean like "add (27 + 15) + (3 + 5)" or something... it is grouping them differently, but not changing the order.

OpenStudy (anonymous):

associative property is \((a+b)+c=a+(b+c)\) which allows you to write without parentheses, but there are no parentheses in the question

OpenStudy (aripotta):